What is the minimum mass of ethylene glycol (C2H6O2) that must be dissolved in 14.5 kg of water to prevent the solution from freezing at −18.6°F? (Assume ideal behavior.)

g?
Answer #1

Sol . As Freezing point of solution = - 18.6 °F = (-18.6 - 32) × 5/9 = - 28.11°C

Freezing point of solvent (water) = 0°C

So , Depression in freezing point = deltaTf = Freezing point of solvent - Freezing point of solution

= 0 - ( - 28.11)  

= 28.11°C

Now , deltaTf = Kf × m

where kf = freezing point depression constant

m = molality of solution

For water , Kf = 1.86 °C/m

So , m = deltaTf / Kf = 28.11 / 1.86 = 15.1129 m

As Molar mass of ethylene glycol = 62.07 g/mol

Mass of water = 14.5 Kg

So,

molality of solution = Mass of ethylene glycol (in g)

/ ( Molar mass of ethylene glycol (in g/mol) × Mass of water (in Kg) )

And , Mass of ethylene glycol ( in g) = 15.1129 × 62.07 × 14.5  

= 13601.8366 g